OpenStudy (anonymous):

what's the horizontal asymptote of (8/4x) +1?

OpenStudy (johnweldon1993):

To find the horizontal asymptote...you find what the equation equals when you let x tend to infinity \[\frac{ 8 }{ 4x } + 1\] What happens when 'x' approaches infinity? what does this equation come out to?

OpenStudy (johnweldon1993):

Hint...what happens to 8/4x when the denominator gets larger and larger? what does it get close it?

OpenStudy (johnweldon1993):

Does it make sense? heres another way of looking at it what does 1/10 =? a small number now what does 1/100 =? an even smaller number now what does 1/1000000000000000 =? a microscopic number As the denominator gets bigger and bigger...the whole thing gets closer and closer to...?